# MMAR: Towards Lossless Multi-Modal Auto-Regressive Probabilistic Modeling

### v2

@@ Abstract
Recent advancements in multi-modal large language models have propelled the development of joint probabilistic models capable of both image understanding and generation. However, we have identified that recent methods inevitably suffer from loss of image information during understanding task, due to either image discretization or diffusion denoising steps. To address this issue, we propose a novel Multi-Modal Auto-Regressive (MMAR) probabilistic modeling framework. Unlike discretization line of method, MMAR takes in continuous-valued image tokens to avoid information loss. Differing from diffusion-based approaches, we disentangle the diffusion process from auto-regressive backbone model by employing a light-weight diffusion head on top each auto-regressed image patch embedding. In this way, when the model transits from image generation to understanding through text generation, the backbone model’s hidden representation of the image is not limited to the last denoising step. To successfully train our method, we also propose a theoretically proven technique that addresses the numerical stability issue and a training strategy that balances the generation and understanding task goals. Through extensive evaluations on 18 image understanding benchmarks, MMAR demonstrates much more superior performance than other joint multi-modal models, matching the method that employs pretrained CLIP vision encoder, meanwhile being able to generate high quality images at the same time. We also showed that our method is scalable with larger data and model size. Code will be available at https://github.com/ydcUstc/MMAR.
@@ 1 Introduction
Over the past few years, extensive research in the field of multimodal intelligence has catalyzed the accelerated advancement of foundational models for both image understanding and image generation. Within the realm of image understanding, multimodal large language models (MLLM), exemplified by LLaVA (Liu et al., 2023a), have exhibited human-like capabilities in open-domain image comprehension. In the domain of image generation, techniques rooted in generative probabilistic models, such as Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) and auto-regressive (AR) models (Chen et al., 2020), have also garnered significant success. Essentially, these two lines of research correspond to modeling the conditional probability, i.e., $P(T|I)$ and $P(I|T)$, where $T$ and $I$ corresponds to text and image, respectively. It’s evident that both types of conditional probabilistic models are subsets of a joint probabilistic model, $P(T,I)$. This brings us to an intriguing question: Could a joint probabilistic model serve as a natural unifying framework for both image understanding and image generation tasks?
@@
Given that the most advanced image understanding (Chen et al., 2024) and generation (Esser et al., 2024) models rely on the language priors $p(T)$ of pre-trained large language models (LLMs), the most straight-forward approach for joint image-text probabilistic modeling is to convert images into discrete tokens similar to text. This way, images are treated as a form of language and integrated into the auto-regressive modeling of text, as seen in models like MARS (He et al., 2024), LlamaGen (Sun et al., 2024a), and Chameleon (Team, 2024). This method leverages the powerful text modeling capabilities of various open-source pre-trained LLMs, along with their highly optimized training and inference frameworks. However, discretization on the continuous image space introduces an information bottleneck, inevitably leading to a loss of image details and reducing the model’s information capacity. This limitation is quantitatively evident in the image understanding performance: existing methods based on image token discretization fall short when compared to the LLaVA model, which utilizes continuous CLIP representations (Xie et al., 2024).While there are theoretically feasible methods to mitigate this information bottleneck, such as significantly increasing the number of tokens or the size of the VQ codebook, these approaches would substantially increase the training difficulty for both LLMs and VQ-VAE models.
@@
Recently, several attempts have been made to address the continuous nature of images by combining image diffusion transformers and text auto-regression within a unified transformer architecture (Zhou et al., 2024; Zhao et al., 2024). Although this approach, exemplified by Transfusion (Zhou et al., 2024) and MonoFormer, demonstrated superior performance compared to discrete token auto-regressive methods for images, it is crucial to acknowledge that the inherent nature of image diffusion prevents this approach from leveraging complete image modeling capability when it comes to image understanding tasks, as shown in fig. 1-(B). The root cause is that image diffusion will encode different parts of image information into different noise levels (Ho et al., 2020). For example, at low noise level, the model tends to solve an image enhancement task, while at high noise level, the model is only suitable for extracting rough image morphology and layout. Intuitively, image-to-text tasks need to take clean images as input to maintain complete visual information. However, due to the nature of diffusion modeling, this will only invoke the hidden representations that are related to the image enhancement task according to its low noise level, resulting in insignificant help for image-to-text tasks. Adding some noise to the input image would help invoke more semantic representations, but at a cost of image information loss. To ensure a full utilization of the image modeling capacity learned by image diffusion, it is necessary to input the images of all noise levels, but with tremendous computational overhead. This dilemma hinders transfusion-like models perform well for visual understanding tasks.
@@
Figure 1: Comparison of different multi-modal probabilistic modeling paradigms. (A): Most AR-based methods adopt discrete image tokens produced by VQ-VAEs. The VQ operation bottlenecks the representation’s information capacity, resulting in information loss. (B): Concurrent transfusion-like methods combining image DiT with text AR. $X^{(i)}$ denotes the full image at noise level $i$. For DiT part, the image information is modeled within the learned representation across various noise level, represented by $h^{(0)},...,h^{(t)}$, where the superscript indicates noise level. However, only $h^{(0)}$ is utilized in image to text tasks. (C): Our MMAR paradigm is VQ-free, thus reducing the information loss related to VQ operation. It auto-regressively models the continuous image tokens $x_{1},...,x_{n}$. The auto-regressive nature guarantees that all the image hidden representations, denoted by $h_{i},i\in[1,n]$, are simultaneously optimized for both $P(I)$ and $P(T|I)$, ensuring complete usage of image modeling capability for visual understanding tasks.
@@
In summary, how to take in and utilize the complete information of the continuous image modality is the major pain point of joint image-text probabilistic modeling. This is the challenge that our work is trying to address. The method proposed in this paper, MMAR, belongs to the image-text joint AR methodology, thus naturally avoids the challenge of fusing diffusion and AR, as mentioned previously. Different from other image-text joint AR methods, our method is built upon continuous image representation that can faithfully reconstruct the original image, rather than the discretized lossy image representation. To model the continuous probability distribution within the auto-regressive paradigm, we refer to MAR (Li et al., 2024b), introducing a simple mlp diffusion sampler to sample continuous image tokens given the transformer backbone output as its conditional information. This technique can also be regarded as disentangling the diffusion process from the auto-regressive backbone in the transfusion-like methods (Zhou et al., 2024; Zhao et al., 2024), which in turn focuses on learning complete visual representation that can guide the entire diffusion sampling process of image tokens. Different from MAR (Li et al., 2024b), we leverage LLMs to learn much more diverse distribution of the large-scale image-text data. In practice, we find that low-precision training strategies for contemporary LLMs introduce non-negligible numerical error term in the diffusion sampler, which substantially impedes the convergence. We carefully analyze the numerical error term and find that the only way to minimize it is to properly parameterize the diffusion model. Therefore, we theoretically derive an optimal diffusion parameterization to solve this issue. Additionally, we find that visual understanding and generation prefer different experimental settings. To balance these two tasks in a unified model, we propose a two-stage training method, which firstly improves visual understanding ability with large, mid quality data, then boost visual generation ability with high quality data. With these innovations, the lossless multi-modal auto-regressive probabilistic modeling is finally achieved theoretically and practically.
@@
  • We identified why previous joint probabilistic models for both image generation and understanding suffer image information loss with a series of experiments.
  • We proposed a novel modeling framework that solves the issue by disentangling diffusion process from auto-regressive backbone, and achieved significant performance improvement for image understanding task.
@@ 2.1 Multi-modal Large Language Models
Since LLMs demonstrated open-domain conversational capabilities, more and more research has be focused on how to introduce visual information into large language models to achieve open-domain visual understanding. Pioneering works such as BLIP-2 (Li et al., 2023b), MiniGPT4 (Zhu et al., 2024) and LLaVA (Liu et al., 2023a) use trainable connector modules such as qformer or mlp to align image representations to the input space of LLM, making open-domain visual question answering possible. Recently, thanks to innovations in network architecture (Dong et al., 2024b; Tong et al., 2024), training strategy (Chen et al., 2024) and the support for dynamic resolution input (Hu et al., 2024; Li et al., 2024a), the visual understanding performance of large multi-modal models have been greatly improved. These works focus on the alignment of image representations to text representations, only achieving $p(T|I)$ without incorporating the image’s own distribution $p(I)$ into the modeling capabilities of the model. Different from these works, our work adds the modeling of $p(I)$ on the basis of these MLLMs to achieve joint image-text probabilistic modeling.
@@ 2.2 Auto-Regressive Image Generative Models
Text-to-image generative models aim at modeling the conditional probability distribution $p(I|c)$, enabling probabilistic sampling of images conditioned on textual or categorical inputs. Auto-regressive methods represent a dominant paradigm in this domain, typically requiring discrete representations for both input and output. For images, this necessitates encoding them into discrete codes using a VQVAE (Esser et al., 2021; Kondratyuk et al., 2024). While recent works demonstrate that auto-regressive methods based on discrete image tokens can generate high-quality images (Sun et al., 2024a), the discretization of image representation acts as an information bottleneck, limiting the modeling accuracy of the image distribution. Recent efforts have shown that auto-regressive probabilistic modeling can be achieved without relying on discrete representations (Tschannen et al., 2023; Li et al., 2024b). For instance, MAR (Li et al., 2024b) replaces traditional logits with diffusion heads, enabling probabilistic sampling of continuous representations within an auto-regressive framework. This paper introduces continuous representation auto-regressive probability modeling to MLLMs, mitigating information loss caused by quantization and achieving theoretically lossless joint image-text probability modeling. In addition, we addressed the difficulty when training with large model and large data which is not presented in the previous works.
@@ 2.3 Unified Visual Understanding and Generation Models
Recently, a series of studies have focused on leveraging a single model to simultaneously address tasks of visual generation and understanding. Early works in this area adopted a modular approach, bridging pre-trained models for visual understanding and visual generation through intermediate representations to achieve combined image understanding and generation. Notable examples include EMU (Sun et al., 2024c, b) and SEED-X (Ge et al., 2024). These works, however, are not considered probabilistic modeling because they aim at modeling the mean value of representations like CLIP or other intermediate representations rather than modeling the true image distribution. This limitation leads to the inadequate image space exploration, and thus hinders the attainment of high-quality generative and understanding performance.
Another line of research adheres to the paradigm of probabilistic modeling (Team, 2024; He et al., 2024; Xie et al., 2024; Wu et al., 2024; Zhou et al., 2024; Zhao et al., 2024). These approaches can be categorized into three types based on whether the image representations are discrete and the modeling method of the image part : (i) Discrete auto-regressive Methods: Examples include Chameleon (Team, 2024), MARS (He et al., 2024), and VILA-U (Wu et al., 2024). These methods discretize image representations and then model images and text tokens using a unified auto-regressive transformer. (ii) Discrete Diffusion Methods: An example is Show-o (Xie et al., 2024). These methods discretize image tokens and model them with text tokens using a unified transformer through a discrete diffusion approach. (iii) Continuous Diffusion Methods: Examples include Transfusion (Xie et al., 2024) and MonoFormer (Zhao et al., 2024). These methods do not discretize image representations but directly employ continuous diffusion transformers to model image tokens along with text tokens using a unified transformer. Our approach differs from the aforementioned three types. It belongs to the continuous auto-regressive method category, which does not require discretizing image representations. Instead, it predicts the continuous distribution of image tokens using an auto-regressive approach and models them alongside text tokens within a unified transformer, as shown in fig.1.
@@ 3.1 Auto-Regressive Modeling with Continuous and Discrete Representations
In the realm of natural language processing (NLP), the sequence $\mathbf{x}$ is solely made of discrete text tokens. As a result, most modern large language models (LLMs) parameterize $p_{\theta}(x_{i}|x_{<i})$ into a categorical distribution, which can be explicitly represented by the softmax activation on a set of logits predicted by a decoder-only transformer (Radford et al., 2019; Brown et al., 2020) $f_{\theta}(\cdot)$ followed by a linear LM head $H_{\theta}(\cdot)$:
@@
In addition to text, our work also aims at modeling the probability of images, which are represented by continuous rather than discrete image tokens. Therefore, a protocol for parameterizing $p_{\theta}(x_{i}|x_{<i})$ of the continuous image tokens is required. Inspired by MAR (Li et al., 2024b), we train a diffusion model to achieve this. The diffusion model takes vector $z_{i}=f_{\theta}(x_{<i})$ as the conditional input, and samples $x_{i}$ by gradually denoising from a randomly sampled gaussian noise. To optimize the diffusion model for continuous image token sampling, a diffusion loss can be applied, which acts as the upper-bound of the negative log likelihood. A typical diffusion loss can be written as follows, which is seen in MAR (Li et al., 2024b):
$L(x_{i})=\mathbb{E}_{x_{i},\epsilon,t}[w_{t}\cdot||\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}]\geq-\log p_{\theta}(x_{i}|x_{<i})+C,$ (3)
where $w_{t}$ is the loss weight that balances the loss for different timesteps, and $\bar{\alpha}_{t}$ indicates the noise schedule of the forward diffusion process. In this way, minimizing the diffusion loss is equivalent to maximizing the log likelihood of image data.
The overall loss for joint image-text probabilistic modeling can be written as follows:
$L=\sum_{i\in I_{img}}L(x_{i})-\sum_{i\in I_{txt}}\log p_{\theta}(x_{i}|x_{<i}),$ (4)

### v3

@@ Abstract
Recent advancements in multi-modal large language models have propelled the development of joint probabilistic models capable of both image understanding and generation. However, we have identified that recent methods suffer from loss of image information during understanding task, due to either image discretization or diffusion denoising steps. To address this issue, we propose a novel Multi-Modal Auto-Regressive (MMAR) probabilistic modeling framework. Unlike discretization line of method, MMAR takes in continuous-valued image tokens to avoid information loss in an efficient way. Differing from diffusion-based approaches, we disentangle the diffusion process from auto-regressive backbone model by employing a light-weight diffusion head on top each auto-regressed image patch embedding. In this way, when the model transits from image generation to understanding through text generation, the backbone model’s hidden representation of the image is not limited to the last denoising step. To successfully train our method, we also propose a theoretically proven technique that addresses the numerical stability issue and a training strategy that balances the generation and understanding task goals. Extensive evaluations on 18 image understanding benchmarks show that MMAR significantly outperforms most of the existing joint multi-modal models, surpassing the method that employs pre-trained CLIP vision encoder. Meanwhile, MMAR is able to generate high quality images. We also show that our method is scalable with larger data and model size. Code will be available at https://github.com/ydcUstc/MMAR.
@@ 1 Introduction
Over the past few years, extensive research in the field of multimodal intelligence has catalyzed the accelerated advancement of foundational models for both image understanding and image generation. Within the realm of image understanding, multimodal large language models (MLLM), exemplified by LLaVA [33], have exhibited human-like capabilities in open-domain image comprehension. In the domain of image generation, techniques rooted in generative probabilistic models, such as Denoising Diffusion Probabilistic Models (DDPM) [21] and auto-regressive (AR) models [6], have also garnered significant success. Essentially, these two lines of research correspond to modeling the conditional probability, i.e., $P(T|I)$ and $P(I|T)$, where $T$ and $I$ corresponds to text and image, respectively. It’s evident that both types of conditional probabilistic models are subsets of a joint probabilistic model, $P(T,I)$. This brings us to an intriguing question: Could a joint probabilistic model serve as a natural unifying framework for both image understanding and image generation tasks?
@@
Given that the most advanced image understanding [8] and generation [14] models rely on the language priors $p(T)$ of pre-trained large language models (LLMs), the most straight-forward approach for joint image-text probabilistic modeling is to convert images into discrete tokens similar to text. This way, images are treated as a form of language and integrated into the auto-regressive modeling of text, as seen in models like MARS [19], LlamaGen [53], and Chameleon [57]. This method leverages the powerful text modeling capabilities of various open-source pre-trained LLMs, along with their highly optimized training and inference frameworks. However, the code-book size strictly restricts the information capacity per token, leading to a loss of image details and reducing the model’s information capacity when the image token number is small. This limitation is quantitatively evident in the image understanding performance: most of the existing methods based on image token discretization [57, 62] fall short when compared to the LLaVA model, which utilizes continuous CLIP representations. Although some approaches can mitigate this bottleneck, such as increasing the number of tokens (4096 tokens per 512x512 image for EMU3 [60]) or increasing the size of the VQ code-book, these approaches would substantially increase the training cost for both LLMs and VQ-VAE models. In contrast, without the restriction of code-book size, continuous image tokenizers can compress images much more efficiently. Recent research [4] even shows that it is possible to compress a 128x128 image patch into a single continuous image token without significant detail loss, resulting in only 16 tokens per 512x512 image. This indicates huge potential of using continuous image tokens instead of discrete ones.
Figure 1: The trend of MMBench tasks’ performance of a transfusion-like model with respect to input image noise levels. (A): The best performance of most tasks are achieved at a non-zero noise level. (B): The object localization performance consistently gets higher as the noise level grows up.
@@
Recent efforts on leveraging continuous image tokens include transfusion [71] and MonoFormer [70]. They combine image diffusion transformers and text auto-regression within a unified transformer architecture, and demonstrates superior performance compared to image discretization approaches. However, when we reproduce and analyze this approach, we find it has difficulty utilizing complete image modeling capability for image understanding tasks when only the clean image is input to the model. As shown in Fig. 1, the best performance of different image understanding tasks are achieved at different non-zero noise levels. Especially, the object localization task even prefers very noisy inputs. This counter-intuitive phenomenon reveals that, diffusion models encode different parts of image information into different noise levels. At low noise level, the model tends to solve an image enhancement task, while at high noise level, the model is only suitable for extracting rough image morphology and layout [21]. Intuitively, image understanding needs to take clean images as input to maintain complete visual information. However, due to the nature of diffusion modeling, this will only invoke the hidden representations that are related to the image enhancement task according to its low noise level, resulting in insignificant help for image understanding. Adding some noise to the input image would help invoke more semantic representations, but at a cost of image information loss. To ensure a full utilization of the image modeling capacity learned by image diffusion, it is necessary to input the images of all noise levels, but with tremendous computational overhead. This dilemma hinders transfusion-like models perform well for visual understanding tasks.
@@
Figure 2: Comparison of different multi-modal probabilistic modeling paradigms. (A): Most AR-based methods adopt discrete image tokens produced by VQ-VAEs. The VQ operation bottlenecks per-tokens information capacity, resulting in information loss or huge token numbers. (B): Transfusion-like methods combining image DiT with text AR. $X^{(i)}$ denotes the full image at noise level $i$. For DiT part, the image information is modeled within the learned representation across various noise level, represented by $h^{(0)},...,h^{(t)}$, where the superscript indicates noise level. However, only $h^{(0)}$ is utilized in image to text tasks. (C): Our MMAR paradigm utilizes continuous image tokens $x_{1},...,x_{n}$, enjoying efficient and lossless token compression. Its auto-regressive nature guarantees that all the image hidden representations, denoted by $h_{i},i\in[1,n]$, are simultaneously optimized for both $P(I)$ and $P(T|I)$, ensuring complete usage of image modeling capability for visual understanding tasks.
@@
In summary, how to efficiently take in and utilize the complete information of the continuous image modality is the major pain point of joint image-text probabilistic modeling. This is the challenge that our work is trying to address. As shown in Fig. 2, the method proposed in this paper, MMAR, belongs to the image-text joint AR methodology, which ensures the complete use of image modeling capability by its auto-regressive nature. Different from other image-text joint AR methods, our method takes in continuous image tokens that can efficiently maintain the complete image information. To model the continuous probability distribution within the auto-regressive paradigm, we refer to MAR [30], introducing a simple mlp diffusion sampler to sample continuous image tokens given the transformer backbone output as its conditional information. This technique can also be regarded as disentangling the diffusion process from the auto-regressive backbone in the transfusion-like methods [71, 70], which in turn focuses on learning complete visual representation that can guide the entire diffusion sampling process of image tokens. Different from MAR [30], we leverage LLMs to learn much more diverse distribution of the large-scale image-text data. In practice, we find that low-precision training strategies for contemporary LLMs introduce non-negligible numerical error term in the diffusion sampler, which substantially impedes the convergence. We carefully analyze the numerical error term and find that a effective way to minimize it is to properly parameterize the diffusion model. Therefore, we theoretically derive an optimal diffusion parameterization to solve this issue. Additionally, visual understanding and generation prefer different experimental settings. To balance these two tasks in a unified model, we propose a two-stage training method, which firstly improves visual understanding ability with large, mid quality data, then boost visual generation ability with high quality data. With these innovations, the lossless multi-modal auto-regressive probabilistic modeling is finally achieved theoretically and practically.
@@
  • We identified why previous joint probabilistic models for both image generation and understanding suffer image information loss.
  • We are the first to combine continuous image representation and discrete text representation into a unified auto-regressive probabilistic modeling framework. Our framework successfully overcomes information loss while keeping high efficiency, leading to significant performance gains on various image understanding tasks.
@@ 2.1 Multi-modal Large Language Models
Since LLMs demonstrated open-domain conversational capabilities, more and more research has been focused on how to introduce visual information into large language models to achieve open-domain visual understanding. Pioneering works such as BLIP-2 [29], MiniGPT4 [72] and LLaVA [33] use trainable connector modules such as qformer or mlp to align image representations to the input space of LLM, making open-domain visual question answering possible. Recently, thanks to innovations in network architecture [10, 58], training strategy [8] and the support for dynamic resolution input [23, 27], the visual understanding performance of large multi-modal models have been greatly improved. These works focus on the alignment of image representations to text representations, only achieving $p(T|I)$ without incorporating the image’s own distribution $p(I)$ into the modeling capabilities of the model. Different from these works, our work adds the modeling of $p(I)$ on the basis of these MLLMs to achieve joint image-text probabilistic modeling.
@@ 2.2 Auto-Regressive Image Generative Models
Text-to-image generative models aim at modeling the conditional probability distribution $p(I|c)$, enabling probabilistic sampling of images conditioned on textual or categorical inputs. Auto-regressive methods represent a dominant paradigm in this domain, typically requiring discrete representations for both input and output. For images, this necessitates encoding them into discrete codes using a VQVAE [13, 25]. While recent works demonstrate that auto-regressive methods based on discrete image tokens can generate high-quality images [53], the discretization of image representation acts as an information bottleneck, limiting the modeling accuracy of the image distribution. Recent efforts have shown that auto-regressive probabilistic modeling can be achieved without relying on discrete representations [59, 30]. For instance, MAR [30] replaces traditional logits with diffusion heads, enabling probabilistic sampling of continuous representations within an auto-regressive framework. This paper introduces continuous representation auto-regressive probability modeling to MLLMs, mitigating information loss caused by quantization and achieving theoretically lossless joint image-text probability modeling. In addition, we addressed the difficulty when training with large model and large data which is not presented in the previous works.
@@ 2.3 Unified Visual Understanding and Generation Models
Recently, a series of studies have focused on leveraging a single model to simultaneously address tasks of visual generation and understanding. Early works in this area adopted a modular approach, bridging pre-trained models for visual understanding and visual generation through intermediate representations to achieve combined image understanding and generation. Notable examples include EMU [55, 54] and SEED-X [16]. These works, however, are not considered probabilistic modeling because they aim at modeling the mean value of representations like CLIP or other intermediate representations rather than modeling the true image distribution. This limitation leads to the inadequate image space exploration, and thus hinders the attainment of high-quality generative and understanding performance.
Another line of research adheres to the paradigm of probabilistic modeling [57, 19, 62, 61, 71, 70]. These approaches can be categorized into three types based on whether the image representations are discrete and the modeling method of the image part : (i) Discrete auto-regressive methods: Examples include Chameleon [57], MARS [19], VILA-U [61], and EMU-3 [60]. These methods discretize image representations and then model images and text tokens using a unified auto-regressive transformer. (ii) Discrete diffusion methods: An example is Show-o [62]. These methods discretize image tokens and model them with text tokens using a unified transformer through a discrete diffusion approach. (iii) Continuous diffusion methods: Examples include Transfusion [71] and MonoFormer [70]. These methods do not discretize image representations but directly employ continuous diffusion transformers to model image tokens along with text tokens using a unified transformer. Our approach differs from the aforementioned three types. It belongs to the continuous auto-regressive method category, which does not require discretizing image representations. Instead, it predicts the continuous distribution of image tokens using an auto-regressive approach and models them alongside text tokens within a unified transformer.
@@ 3.1 Auto-Regressive Modeling with Continuous and Discrete Representations
In the realm of natural language processing (NLP), the sequence $\mathbf{x}$ is solely made of discrete text tokens. As a result, most modern large language models (LLMs) parameterize $p_{\theta}(x_{i}|x_{<i})$ into a categorical distribution, which can be explicitly represented by the softmax activation on a set of logits predicted by a decoder-only transformer [44, 1] $f_{\theta}(\cdot)$ followed by a linear LM head $H_{\theta}(\cdot)$:
@@
In addition to text, our work also aims at modeling the probability of images, which are represented by continuous rather than discrete image tokens. Therefore, a protocol for parameterizing $p_{\theta}(x_{i}|x_{<i})$ of the continuous image tokens is required. Inspired by MAR [30], we train a diffusion model to achieve this. The diffusion model takes vector $z_{i}=f_{\theta}(x_{<i})$ as the conditional input, and samples $x_{i}$ by gradually denoising from a randomly sampled gaussian noise. To optimize the diffusion model for continuous image token sampling, a diffusion loss can be applied, which acts as the upper-bound of the negative log likelihood. A typical diffusion loss can be written as follows, which is seen in MAR [30]:
$L_{i}$ $=\mathbb{E}_{x_{i},\epsilon,t}[w_{t}\cdot||\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}]$
$\geq-\log p_{\theta}(x_{i}|x_{<i})+C,$ (3)
where $w_{t}$ is the loss weight that balances the loss for different timesteps, and $\bar{\alpha}_{t}$ indicates the noise schedule of the forward diffusion process. In this way, minimizing the diffusion loss is equivalent to maximizing the log likelihood of image data conditioned on its preceding text guidance.
The overall loss for joint image-text probabilistic modeling can be written as follows, which is an upper-bound of the negative log likelihood of the multi-modal data $\mathbf{x}$:
$L=\sum_{i\in I_{img}}L_{i}-\sum_{i\in I_{txt}}\log p_{\theta}(x_{i}|x_{<i})\geq-\log p_{\theta}(\mathbf{x})+C,$ (4)